3,9,15, 6,17, 16,32,24,12, 30, 35,12,26,21,25 Using StatKey or other technology, find the following values for the above data Click here to access StatKey (a) The mean and the standard deviation Round your answers to one decimal place. mean standard deviation (b) The five number summary. Enter exact answers. The five number summary is
Answer #1

a)

Mean: 〉 Xi/n 283/15-18.86667

X Mean (X-Mean)^2
3 18.86667 251.7511
9 18.86667 97.35111
15 18.86667 14.95111
6 18.86667 165.5511
17 18.86667 3.484444
16 18.86667 8.217778
32 18.86667 172.4844
24 18.86667 26.35111
12 18.86667 47.15111
30 18.86667 123.9511
35 18.86667 260.2844
12 18.86667 47.15111
26 18.86667 50.88444
21 18.86667 4.551111
25 18.86667 37.61778
Sum 1311.733

Standard deviation: = V131 1733/15= 9.35141

b)Five number summary:

Minimum: 3

Q1: (n+1)/4 th observation= (15+1)/4 =4th observation= 12 (Incremental order)

Median: (n+1)/2 th observation= (15+1)/2= 8th observation= 17

Q3:3*(n+1)/4 th observation= 3*(15+1)/4 =3*4=12 th observation= 26

Maximum: 35